hello my name is anthony and who ever reponds back to this it would be great cuz i am have trouble with algerbra equations with varibles on each sides and so i am new at this and so if you could tell me the steps and explain to me the best way you can that would be great


5x-3-=3x+7

Answer by rfer(16322)

5x-3=3x+7
subtract 3x from both sides
5x-3x-3=3x-3x+7
2x-3=7
add 3 to both sides
2x-3+3=7+3
2x=10
divide by 2
x=5